OpenStudy (anonymous):

. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + 4y = 1 -6x + 6y = -3

OpenStudy (anonymous):

Can you multiply the first equation by 2:

OpenStudy (anonymous):

2(3x + 4y = 1) -6x + 6y = -3 6x + 8y = 2 -6x + 6y = -3 14y = -1 y = -1/14 3x + 4(-1/14) = 1 3x + -2/7 = 1 3x = 9/7 x = 3/7

OpenStudy (anonymous):

thank you, i appreciate it

OpenStudy (anonymous):

You're welcome